a die is tossed 120 times if number of 3 appears is recorded what is the mean and variance and standard deviation of the experiment
Answer by ewatrrr(24785)


p(3) = 1/6
mean: 1/6(120) = 20
variance 20(5/6) = 100/6
standard deviation of the experiment: sqrt(100/6)